# Tensegrity Robot Proprioceptive State Estimation with Geometric Constraints

### v1

@@ Abstract
Tensegrity robots, characterized by a synergistic assembly of rigid rods and elastic cables, form robust structures that are resistant to impacts. However, this design introduces complexities in kinematics and dynamics, complicating control and state estimation. This work presents a novel proprioceptive state estimator for tensegrity robots. The estimator initially uses the geometric constraints of 3-bar prism tensegrity structures, combined with IMU and motor encoder measurements, to reconstruct the robot’s shape and orientation. It then employs a contact-aided invariant extended Kalman filter with forward kinematics to estimate the global position and orientation of the tensegrity robot. The state estimator’s accuracy is assessed against ground truth data in both simulated environments and real-world tensegrity robot applications. It achieves an average drift percentage of 4.2%, comparable to the state estimation performance of traditional rigid robots. This state estimator advances the state of the art in tensegrity robot state estimation and has the potential to run in real-time using onboard sensors, paving the way for full autonomy of tensegrity robots in unstructured environments.
@@ I Introduction
Tensegrity robots are composed of rigid rods suspended by a network of cables or tendons. This tensional integrity structural design enhances the robot’s resilience, allowing it to absorb impacts and undergo deformation without losing functionality[1, 2]. By modulating the cable length, the robots can morph their shape on the fly, enabling them to maneuver through constrained environments and traverse unstructured terrains. Most existing research on tensegrity robots has predominantly focused on their design[3, 4, 5, 6], control[7, 8, 9], and dynamics[7, 2], with relatively limited attention given to the challenge of state estimation. The inherent flexibility and deformable structure of the design, combined with their unique rolling maneuver, introduce significant challenges in accurately reconstructing the robots’ shape and estimating their pose.
Figure 1: Tensegrity robot state estimation experiment setup, the robot is teleoperated to roll on a flat ground.
Autonomous robots with motion planning and feedback control capabilities rely extensively on accurate state estimation, i.e., position, orientation, and velocity. Exteroceptive sensors, e.g., cameras, LiDARs, are commonly used to provide perceptual data for global localization and navigation[10, 11]. However, these sensors typically operate at low frequencies(10-30Hz) and are susceptible to featureless scenarios[12, 13] and motion blurs[14]. In contrast, proprioceptive sensors, e.g., IMUs and joint encoders, are more robust against external disturbances such as illumination changes and can run at high frequencies (50-1000Hz).
@@
In this paper, we developed a proprioceptive state estimator for the tensegrity robot using an invariant extended Kalman filter (InEKF). The proposed method estimates the robot shape in the body frame as kinematics information, further integrating the inertial measurements to estimate the robot’s position and orientation. The method is evaluated quantitatively in both Mujoco simulation and real-world settings, where the tensegrity robot is locomoted using pre-computed gaits. The estimated pose is compared to ground truth obtained from the simulation or a motion capture system, and achieves an average of 4.2% final drift percentage, comparable to wheeled robots at 3.8%, legged robots at 1.65 % and full-size vehicles at 3.18% as reported in [13].
@@
    We propose the first proprioceptive InEKF state estimator that can estimate both the shape and pose of tensegrity robot system, introducing a measurement model tailored to the kinematics of tensegrity robots.
@@
    We incorporate the geometric properties of the 3-bar tensegrity robot as constraints in the robot shape reconstruction process using constrained optimization.
@@
    We validated the effectiveness of the proposed method through both simulation and hardware experiments conducted with the 3-bar tensegrity robot rolling on flat terrain.
@@ II Related Work
Proprioceptive state estimation problems for traditional rigid robots are well-studied. For legged robots, the initial proprioceptive state estimation methods were based on leg kinematics, known as legged odometry (LO)[22, 23]. Legged odometry assumes the contact foot is static in the world frame, and the robot’s state is estimated using the motion of the legs, including position and orientation, through forward kinematics based on joint angles and foot contacts. More recent work has focused on incorporating IMU data with legged odometry(LIO) using Kalman Filter (KF)[24]. Bloesch et al. [25] proposed estimating robot pose, velocity, and IMU biases using body IMU, joint encoders, and foot contact sensors. Hartley et al.[26] designed a contact-aided Invariant Extended Kalman Filter (InEKF) for legged robot state estimation. The IMU strapdown model was used for propagation, while foot contact positions were augmented in the state during the correction step using forward kinematics. The contact-aided InEKF improved orientation estimation convergence and overall performance on the bipedal Cassie robot.
@@
Wheeled robots and autonomous vehicles, on the other hand, utilize data from IMU and wheel encoders to provide robust state estimates, even in the presence of collisions or slippage events [27, 28]. Legged and wheeled robots benefit from deterministic kinematics, where control inputs are directly mapped to body velocity. However, tensegrity robots pose a unique challenge, as their control algorithms and kinematics depend on continuously changing shapes. This makes it crucial to estimate both the robot’s shape and pose in real-time.
@@
For shape reconstruction, which involves estimating the endcap positions within the body frame using onboard sensors, several approaches have been explored. Elastic stretch sensors, as demonstrated in [29, 18], can estimate the distance between two cable-connected endcaps. In [5], motor encoders provide accurate cable length measurements for the nylon cables connecting the endcaps. IMUs installed on the structure can estimate the robot’s orientation [30]. Li et al.[20] proposed a new shape recognition method using elastic cable length sensors and recurrent neural network (RNN), although this approach overlooks the tensegrity robot’s geometrical properties. Booth et al. [31] developed a sensor skin with conductive capacitive strain sensors to measure the distance between adjacent endcaps and used optimization to determine the endcap positions within the local reference frame.
@@
For pose estimation, Moldagalieva et al. [32] proposed a vision-based method for arm-like tensegrity robots using fiducial markers assuming that the markers remain visible to overhead cameras. Previous work [33] combined proprioceptive and exteroceptive sensors, using an Unscented Kalman Filter (UKF) to estimate global rods poses by fusing data from IMU and (Ultra-Wideband) UWB positioning sensor placed at the endcaps of the rods. More recently, Lu et al. [18] track the 6-DoF pose of each rod using overhead depth and RGB images, combined with cable length measurements from onboard stretch sensors. Their method uses the correspondence between depth and RGB images to estimate rod poses during the transition step. In the correction step, joint-constrained optimization is performed with rod pose estimates and cable length measurements, to refine estimates of the global rod poses.
@@ III-A Robot definition
As shown in Fig. 2, the 3-bar triangular prism tensegrity robot consists of three identical rigid rods with endcaps on each end. The endcaps on the same side denoted as $\{s_{0},s_{2},s_{4}\}$ and $\{s_{1},s_{3},s_{5}\}$ in Fig. 2—are connected by three length controlled side cables. The endcaps on one end of each rod are also connected to the opposite endcaps of adjacent rods by three cross cables arranged in a counterclockwise configuration, mimicking elastic springs. The shape-morphing capability is leveraged to enable rolling locomotion by shifting the robot’s center of mass.
Unlike other robots, the tensegrity robot is rotational symmetric, complicating the definition of a fixed robot base—essential for establishing the mathematical representation and kinematics of the robots. To address this, we designate the rod $r_{01}$ in Fig. 2 as the base rod. The origin of the robot’s body frame is defined at the position of the IMU, installed near the center of this base rod, denoted as $c_{01}$, with an offset of $d_{\text{offset}}$ towards $s_{0}$. The z-axis is aligned along the axial direction of the base rod. This specific frame definition is critical for developing the kinematic model and accurately describing the robot’s motion.
@@ III-B Problem definition
This work aims to track the pose of the 3-bar tensegrity robot and the robot shape in body frame. i.e.,
  • Tensegrity robot body pose in world frame at time $t$: $\mathbf{x_{t}}\in\mathrm{SE(3)}$,
@@
  • IMU acceleration and angular velocity measurements at time $t$: $\tilde{a}_{t}$, $\tilde{\omega}_{t}$,
@@
  • $L_{\text{rod}}$ is a constant that represents the rigid bar’s length,
@@ IV Methodology
Our assumptions are as follows:
  • All nine cable lengths can be accurately measured for robot shape reconstruction. This can be achieved by using stretch sensors [19, 18] or motor encoders [5].
  • Contact events between the robot endcap and ground can be reliably detected.
  • The endcap-ground contact is modeled as point contact and remains static in the single contact period.
@@ IV-A State Representation on Lie Group
In robot state estimations, our focus is on the robot’s orientation, velocity, and position relative to a fixed world frame. We denote the rotational matrix $\mathbf{R}_{t}\in\mathrm{SO(3)}$ to represent the transformation from the body frame to the world frame and $\mathbf{v}_{t}\in\mathbb{R}^{3}$ and $\mathbf{p}_{t}\in\mathbb{R}^{3}$ as the body velocity and position in the world frame, respectively. In InEKF framework[26, 27], the state is defined in direct isometries group $\mathbf{x_{t}}\in\mathrm{SE_{l+2}(3)}$ [34, 35, 36, 13]:
@@
where $\mathbf{d}^{l}_{t}\in\mathbb{R}^{3}$ is added as augmented states for foot contact position in the world frame if valid static contact exists.
@@ IV-B IMU Propagation
The IMU measures acceleration $a_{t}\in\mathbb{R}^{3}$ and angular velocity $\omega_{t}\in\mathbb{R}^{3}$. We model the IMU measurements as being corrupted by additive multivariate Gaussian noise:
@@
We assume that the endcap contact position in the world frame remains static throughout the contact period. However, in order to compensate for slippage and other uncertainties, we model the contact velocity as Gaussian white noise $w_{t}^{d}\sim\mathcal{N}(0,\Sigma^{d})$. The unbiased continuous system dynamics can then be expressed as:
@@
where, $\mathbf{w}_{t}:=\mathrm{vec}(w^{g}_{t},w^{a}_{t},\mathbf{0}_{3,1},h_{R}(\mathrm{L}_{t},\tilde{a}_{t})w_{t}^{d})$ and $u_{t}:=[(\tilde{\omega}_{t})^{\mathsf{T}},(\tilde{a}_{t})^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$. We can verify that the dynamic function $f_{u_{t}}(\cdot)$ satisfies the group affine condition and the right invariant error is defined as: $\eta_{t}^{r}=\bar{X}_{t}X_{t}^{-1}=\left(\bar{X}_{t}L\right)\left(X_{t}L\right)^{-1}$, where, $\bar{X_{t}}$ and $X_{t}$ represent two trajectories in the group $\mathcal{G}$, and $L\in\mathcal{G}$ is an arbitrary element of the group.
@@ IV-C2 Correction Model
$+\begin{bmatrix}J_{p}(\mathrm{L}_{t},\tilde{a}_{t})w^{a}_{t_{k}}\\ 0\\ 0\end{bmatrix},$
@@
Finally, the RI-EKF contact correction can be calculated using the following:
@@
where the Kalman gain $L_{t_{k}}$ and innovation covariance $S_{t_{k}}$,
@@ V-A Optimization Formulation
$\textbf{q}_{t}^{2}-\textbf{q}_{t}^{4},\textbf{q}_{t}^{5}-\textbf{q}_{t}^{1}\right>>0$ (15j)
$\textbf{q}_{t}^{0}-\textbf{q}_{t}^{2},\textbf{q}_{t}^{3}-\textbf{q}_{t}^{5}\right>>0$ (15k)
$\textbf{q}_{t}^{4}-\textbf{q}_{t}^{0},\textbf{q}_{t}^{1}-\textbf{q}_{t}^{3}\right>>0$ (15l)
@@ V-B Optimization variable & objectives
$\mathbf{q^{i}_{t}}$ represents the body frame position of endcap $s_{i}$ at time $t$. The set of positions to be optimized is $\mathrm{Q}_{t}=\left\{\mathbf{q}^{i}_{t}\in\mathbb{R}^{3}|i\in\{0,1,2,3,4,5\}\right\}$. The flattened optimization variable $\mathbf{q}_{t}:={[{\textbf{q}_{t}^{0}}^{\mathsf{T}},{\textbf{q}_{t}^{1}}^{\mathsf{T}},{\textbf{q}_{t}^{2}}^{\mathsf{T}},{\textbf{q}_{t}^{3}}^{\mathsf{T}},{\textbf{q}_{t}^{4}}^{\mathsf{T}},{\textbf{q}_{t}^{5}}^{\mathsf{T}}]}^{\mathsf{T}}$.
@@
V-C Constraints
The equality and inequality constraints of the optimization problem are the following:
Endcap arrangement constraints: Eqn. (15b)-(15c) represents the fixed coordinates of $s_{0}$, $s_{1}$ in the body frame as demonstrated in Fig. 2. Eqn. (15d)-(15e) represents the third element(z-axis) of $\mathbf{q}^{2},\mathbf{q}^{4}$ and $\textbf{q}^{3},\textbf{q}^{5}$ are $>0$ and $<0$ , respectively, i.e. $s_{0},s_{2},s_{3}$ are in the upper-half-plane, $s_{1},s_{3},s_{5}$ are in the lower-half-plane of the body frame, as shown in Fig. 2.
Rod length constraints: Eqn. (15f) are constraints between endcaps. Rods in tensegrity robots are rigid elements. The distance between two rod-connected endcaps should be a constant length $L_{\text{rod}}$ at any time.
Geometric Constraints: Eqn. (15g)-(15i) are geometric constraints to restrict the rods from crossing each other along the time. The $\mathbf{c}_{t}^{ij}$ represents the coordinates of the center of the rigid bar $r_{ij}$. Therefore, $\mathbf{c}_{t}^{23}-\mathbf{c}_{t}^{01}$ represents the vector from the center point of the rod $r_{01}$, to the center of rod $r_{23}$. Then, the chained center vector’s cross-product should be pointing to the axial axis shown in Fig.5 instead of pointing to the opposite or radial direction. So, the dot product of the cross-product vector and the axial vector will remain positive.
Chirality Constraints: Eqn. (15j)-(15l) are chirality constraints. The tensegrity structures are twisted 3-prism, which have two twist directions as shown in Fig. 4. The tensegrity robot in simulation and the world is the same as the right configuration in Fig. 4, whose angles between two corresponding vectors(e.g., $\textbf{q}^{24},\textbf{q}^{15}$) on the other side are acute which satisfy the constraints.

### v2

@@ Abstract
Tensegrity robots, characterized by a synergistic assembly of rigid rods and elastic cables, form robust structures that are resistant to impacts. However, this design introduces complexities in kinematics and dynamics, complicating control and state estimation. This work presents a novel proprioceptive state estimator for tensegrity robots. The estimator initially uses the geometric constraints of 3-bar prism tensegrity structures, combined with IMU and motor encoder measurements, to reconstruct the robot’s shape and orientation. It then employs a contact-aided invariant extended Kalman filter with forward kinematics to estimate the global position and orientation of the tensegrity robot. The state estimator’s accuracy is assessed against ground truth data in both simulated environments and real-world tensegrity robot applications. It achieves an average drift percentage of 4.2%, comparable to the state estimation performance of traditional rigid robots. This state estimator advances the state-of-the-art in tensegrity robot state estimation and has the potential to run in real-time using onboard sensors, paving the way for full autonomy of tensegrity robots in unstructured environments.
Index Terms:
Modeling, Control, and Learning for Soft Robots, Localization, Sensor Fusion
@@ I Introduction
Tensegrity robots are composed of rigid rods suspended by a network of cables or tendons. This tensional integrity structural design enhances the robot’s resilience, allowing it to absorb impacts[1, 2], such as surviving drops from height. Furthermore, it provides shape-morphing capabilities that improve locomotion in challenging environments (e.g., navigating through confined spaces) and enables the structure to collapse into smaller volumes, making transportation and deployment more efficient. Most existing research on tensegrity robots has predominantly focused on their design[3, 4, 5, 6], control[7, 8, 9], and dynamics[7, 2], with relatively limited attention given to the challenge of state estimation. The inherent flexibility and deformable structure of the design, combined with their unique rolling maneuver, introduce significant challenges in accurately reconstructing the robots’ shape and estimating their pose.
Figure 1: Tensegrity robot state estimation experiment setup, the robot is teleoperated to roll on flat ground.
Autonomous robots with motion planning and feedback control capabilities rely extensively on accurate state estimation, i.e., position, orientation, and velocity. Exteroceptive sensors, e.g., cameras and LiDARs, are commonly used to provide perceptual data for global localization and navigation[10, 11]. However, these sensors typically operate at low frequencies (10-30Hz) and are susceptible to illumination change, featureless scenarios[12, 13], and motion blurs[14]. Proprioceptive sensors, such as IMUs and joint encoders, offer real-time high-frequency (50-1000Hz) measurements, which can capture highly dynamic motion, such as rolling transitions for tensegrity robots. These sensors are also essential for high-frequency feedback controllers and state estimators and offer greater reliability as they operate independently of environmental conditions, ensuring resilience against external disturbances.
@@
In this paper, we present a proprioceptive state estimator for the tensegrity robot using an invariant extended Kalman filter (InEKF). Our approach first estimates the robot’s shape in the body frame as kinematics information and subsequently integrates the inertial measurements to estimate the robot’s position and orientation. The proposed method is evaluated quantitatively in both Mujoco simulation and real-world settings, where the tensegrity robot is locomoted using pre-computed gaits. The estimated poses are compared against ground truth data obtained from the simulation or a motion capture system, achieving an average of 4.2% final drift percentage, which is comparable to the reported performance metrics for wheeled robots (3.8%), legged robots (1.65 %) and full-size vehicles (3.18%) as noted in [13].
@@
    We introduce the first proprioceptive InEKF state estimator capable of estimating both the shape and pose of the tensegrity robot, featuring a measurement model tailored to the kinematics of tensegrity robots.
@@
    We incorporate the geometric properties of the 3-bar tensegrity robot as constraints in the shape reconstruction process via constrained optimization.
@@
    We validate the proposed method through extensive simulation and real-world experiments using the 3-bar tensegrity robot performing rolling maneuvers over varied terrains.
@@ II Related Work
Proprioceptive state estimation problems for traditional rigid robots are well-studied. For legged robots, the initial proprioceptive state estimation methods were based on leg kinematics, known as legged odometry (LO)[22, 23]. Legged odometry assumes the foot contact position is static in the world frame, and the robot’s state is estimated using the motion of the legs, including position and orientation, through forward kinematics based on joint angles and foot contacts. More recent work has focused on incorporating IMU data with legged odometry (LIO) using Kalman Filter (KF)[24]. Bloesch et al. [25] proposed estimating robot pose, velocity, and IMU biases using body IMU, joint encoders, and foot contact sensors. Hartley et al. [26] designed a contact-aided Invariant Extended Kalman Filter (InEKF) for legged robot state estimation. The IMU strapdown model was used for propagation, while foot contact positions were augmented in the state during the correction step using forward kinematics. The contact-aided InEKF improved orientation estimation convergence and overall performance on the bipedal Cassie robot.
@@
Wheeled robots and autonomous vehicles, on the other hand, utilize data from IMU and wheel encoders to deliver robust state estimates even under collisions or slippage conditions [27, 28]. Legged and wheeled robots benefit from deterministic kinematics, where control inputs are directly mapped to body velocity. However, tensegrity robots pose a unique challenge, as their kinematics are inherently linked to continuously evolving structural configurations during movement. Consequently, accurately estimating the robot’s shape is crucial to generating the substitute forward kinematics, thereby enabling the application of an InEKF state estimation framework as employed in legged robots.
@@
For shape reconstruction, which involves estimating the endcap positions within the body frame using onboard sensors, various approaches have been investigated. Elastic stretch sensors, as demonstrated in [29, 18], are employed to measure the distance between two cable-connected endcaps. In [5], motor encoders provide accurate cable length measurements for the nylon cables connecting the endcaps. IMUs installed on the structure can estimate the robot’s orientation [30]. Li et al. [20] proposed a new shape recognition method using elastic cable length sensors and recurrent neural network (RNN), although this approach overlooks the tensegrity robot’s geometrical properties. Booth et al. [31] developed a sensor skin with conductive capacitive strain sensors to measure the distance between adjacent endcaps and used optimization to determine the endcap positions within the local reference frame.
@@
For pose estimation, Moldagalieva et al. [32] proposed a vision-based method for arm-like tensegrity robots using fiducial markers, assuming that the markers remain visible to overhead cameras. Previous work [33] combined proprioceptive and exteroceptive sensors, using an Unscented Kalman Filter (UKF) to estimate rods poses by fusing data from IMU and (Ultra-Wideband) UWB positioning sensor placed at the endcaps of the rods. More recently, Lu et al. [18] track the 6-DoF pose of each rod using overhead RGBD images combined with cable length measurements from onboard stretch sensors. Their method uses the correspondence between depth and RGB images to estimate rod poses during the transition step. In the correction step, joint-constrained optimization is performed with rod pose estimates and cable length measurements to refine estimates of the global rod poses.
@@ III-A Robot definition
As shown in Fig. 2, the 3-bar triangular prism tensegrity robot consists of three identical rigid rods with endcaps on each end. The endcaps on the same side denoted as $\{s_{0},s_{2},s_{4}\}$ and $\{s_{1},s_{3},s_{5}\}$ in Fig. 2—are interconnected by three length controlled side cables. The endcaps on one end of each rod are also connected to the opposite endcaps of adjacent rods by three cross cables arranged in a counterclockwise configuration, thereby emulating the behavior of elastic springs. This shape-morphing capability is exploited to facilitate rolling locomotion by shifting the robot’s center of mass.
Unlike other robots, the tensegrity robot is rotationally symmetric, complicating the definition of a fixed robot base—essential for establishing the mathematical representation and kinematics of the robots. To address this, we designate the rod $r_{01}$ in Fig. 2 as the base rod. The origin of the robot’s body frame is defined at the position of the IMU, installed near the center of this base rod, denoted as $c_{01}$, with an offset of $d_{\text{offset}}$ towards $s_{0}$. The z-axis is aligned along the axial direction of the base rod. This specific frame definition is critical for developing the kinematic model and accurately describing the robot’s motion.
@@ III-B Problem definition
This work aims to track the pose of the 3-bar tensegrity robot and the robot shape in the body frame. i.e.,
  • Tensegrity robot body pose in world frame: $\mathbf{x_{t}}\in\mathrm{SE(3)}$.
@@
  • IMU acceleration and angular velocity measurements: $\tilde{a}_{t}$, $\tilde{\omega}_{t}$.
@@
  • $L_{\text{rod}}$ is a constant that represents the rigid bar’s length.
@@ IV Methodology
Our assumptions are: a) all nine cable lengths can be accurately measured for robot shape reconstruction, which can be achieved by using stretch sensors [19, 18] or motor encoders [5]; b) contact events between the robot endcap and ground can be reliably detected; and c) the endcap-ground contact is modeled as point contact and remains static during the single contact period.
@@ IV-A State Representation on Lie Group
In robot state estimations, we focus on the robot’s orientation, velocity, and position relative to a fixed world frame. We denote the rotational matrix $\mathbf{R}_{t}\in\mathrm{SO(3)}$ to represent the robot orientation and $\mathbf{v}_{t}\in\mathbb{R}^{3}$ and $\mathbf{p}_{t}\in\mathbb{R}^{3}$ as the body velocity and position in the world frame. In InEKF framework[26, 27], the state is defined in direct isometries group $\mathbf{x_{t}}\in\mathrm{SE_{l+2}(3)}$ [34, 35, 36, 13]
@@
where $\mathbf{d}^{l}_{t}\in\mathbb{R}^{3}$ is added as augmented states for foot contact position in the world frame if valid static contact occurs.
@@ IV-B IMU Propagation
The IMU measures acceleration $a_{t}\in\mathbb{R}^{3}$ and angular velocity $\omega_{t}\in\mathbb{R}^{3}$. We model the IMU measurements as being corrupted by additive multivariate Gaussian noise with biases denoted as $b_{t}^{a},b_{t}^{g}\in\mathbb{R}^{3}$:
@@
We assume that the endcap contact position in the world frame remains static throughout the contact period. However, to compensate for slippage and other uncertainties, we model the contact velocity as Gaussian white noise $w_{t}^{d}\sim\mathcal{N}(0,\Sigma^{d})$. The unbiased continuous system dynamics can then be expressed as:
@@
where, $\mathbf{w}_{t}:=\mathrm{vec}(w^{g}_{t},w^{a}_{t},\mathbf{0}_{3,1},h_{R}(\mathrm{L}_{t},\tilde{a}_{t})w_{t}^{d})$ and $u_{t}:=[(\tilde{\omega}_{t})^{\mathsf{T}},(\tilde{a}_{t})^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$. We can verify that the dynamic function $f_{u_{t}}(\cdot)$ satisfies the group affine condition, and the right invariant error is defined as: $\eta_{t}^{r}=\bar{X}_{t}X_{t}^{-1}=\left(\bar{X}_{t}L\right)\left(X_{t}L\right)^{-1}$, where, $\bar{X_{t}}$ and $X_{t}$ represent two trajectories in the group $\mathcal{G}$, and $L\in\mathcal{G}$ is an arbitrary element of the group.
@@ IV-C2 Correction Model
$+\begin{bmatrix}J_{p}(\mathrm{L}_{t},\tilde{a}_{t})w^{a}_{t_{k}}\\ 0\\ 0\\ {\color[rgb]{0,0,0}0}\end{bmatrix},$
@@
Finally, the RI-EKF contact correction can be calculated using
@@
where the Kalman gain $L_{t_{k}}$ and innovation covariance $S_{t_{k}}$ are
@@ V-A Optimization Formulation
The variable $\mathbf{q}_{t}^{i,j}$, where $j\in{1,2,3}$, denotes the $j$-th component of the endcap position vector $\mathbf{q}_{t}^{i}$. Furthermore, $\mathbf{c}_{t}^{ij}$ represents the coordinates of the center of the rigid bar $r_{ij}$.
@@
$\mathbf{q}_{t}^{2}-\mathbf{q}_{t}^{4},\mathbf{q}_{t}^{5}-\mathbf{q}_{t}^{1}\right>>0$ (15j)
$\mathbf{q}_{t}^{0}-\mathbf{q}_{t}^{2},\mathbf{q}_{t}^{3}-\mathbf{q}_{t}^{5}\right>>0$ (15k)
$\mathbf{q}_{t}^{4}-\mathbf{q}_{t}^{0},\mathbf{q}_{t}^{1}-\mathbf{q}_{t}^{3}\right>>0$ (15l)
@@ V-B Optimization variable & objectives
$\mathbf{q^{i}_{t}}$ represents the body frame position of endcap $s_{i}$ at time $t$. The set of positions to be optimized is $\mathrm{Q}_{t}=\left\{\mathbf{q}^{i}_{t}\in\mathbb{R}^{3}|i\in\{0,1,2,3,4,5\}\right\}$. The optimization variable $\mathbf{q}_{t}:={[{\mathbf{q}_{t}^{0}}^{\mathsf{T}},{\mathbf{q}_{t}^{1}}^{\mathsf{T}},{\mathbf{q}_{t}^{2}}^{\mathsf{T}},{\mathbf{q}_{t}^{3}}^{\mathsf{T}},{\mathbf{q}_{t}^{4}}^{\mathsf{T}},{\mathbf{q}_{t}^{5}}^{\mathsf{T}}]}^{\mathsf{T}}$.
@@
V-C Optimization constraints
The equality and inequality constraints of the optimization problem are as follows:
Endcap arrangement constraints: Eqn. (15b)-(15c) represent the fixed coordinates of $s_{0}$, $s_{1}$ in the body frame as shown in Fig. 2. Eqn. (15d)-(15e) represent the third element (z-axis) of $\mathbf{q}^{2},\mathbf{q}^{4}$ and $\mathbf{q}^{3},\mathbf{q}^{5}$, which are $>0$ and $<0$ , respectively, i.e., $s_{0},s_{2},s_{3}$ are in the upper-half-plane, while $s_{1},s_{3},s_{5}$ are in the lower-half-plane of the body frame.
Rod length constraints: Eqn. (15f) defines constraints between endcaps. Since the rods in tensegrity robots are rigid, the distance between any two endcaps connected by a rod must remain invariant, equaling $L_{\text{rod}}$ at all times.
Geometric Constraints: Eqn. (15g)-(15i) are geometric constraints to prevent the rods from crossing each other over time. $\mathbf{c}_{t}^{23}-\mathbf{c}_{t}^{01}$ represents the vector from the center point of the rod $r_{01}$, to the center of rod $r_{23}$. The cross-product of the chained center vectors should point along the axial axis, as shown in Fig. 5, rather than in the opposite or radial direction. Thus, the dot product of the cross-product vector and the axial vector remains positive.
Chirality Constraints: Eqn. (15j)-(15l) are chirality constraints. The tensegrity structure forms a twisted 3-prism, which can have two twist directions as shown in Fig. 4. The tensegrity robot in both the simulation and the real world follows the right configuration in Fig. 4, where the angles between two corresponding vectors (e.g., $\mathbf{q}^{24},\mathbf{q}^{15}$) on the other side are acute, satisfying the constraints.